A biochemist idenfies a toxin small molecule secreted by a specific type of bacteria. She performs an analysis and determines that a sample contains 0.288 g H, 1.71 g C, and 0.999 g N. She also determines that the molar mass of this toxin is between 120 and 130 g/mol. Using this information, determine the molecular formula of the compound. (C H N)
Answer #1

Answer:

Lets calculate moles of C, H and N in the given compound

Moles of C = 1.71 g/ (molar mass of C) = 1.71 g / 12.0 g/mol = 0.1425 moles

Moles of H = 0.288 g / molar mass of H = 0.288 / 1 g/mol = 0.288 moles

Moles of N = 0.999 g molar mass of N = 0.999 g/ 14.0 g/mol = 0.0713 moles

Molar ratio of C to H to N is = Moles of C : Moles of H : Moles of N = 0.1425 : 0.288 : 0.0713

Simplifying above ratio gives = Moles of C : Moles of H : Moles of N = 2:4:1 (approximately)

So the empirical formula of compound will be C2H4N

But molar mass is between 120 g/mol to 130 g/mol

Mass of C2H4N unit is = (12.0 x 2 + 1.0 x 4 + 14.0 x 1)g/mol = 42 g/mol

So 42 x 3 = 126 and 126 is between 120 and 130

So final formula have 3 units of C2H4N

So final formula (Molecular formula) is C6H12N3